# Omnigenous umbilic stellarators

### v1

@@ 5 Umbilic coil design for a finite-$\beta$ equilibrium
As we have seen in section 4.1, creating modular coils for vacuum umbilic equilibria is a difficult task due to the sharp curvature regions. Therefore, in this section, we present an alternate strategy that uses umbilic coils to create the right boundary shape for a finite-$\beta$ equilibrium. Instead of modifying the modular coils, we can use an umbilic coil, a low-current helical coil which can be used to create an umbilic edge. In the following sections, we will discuss two cases, categorized by the direction of the current in the umbilic coils with respect to the toroidal plasma current.
5.1 Umbilic coil current opposite to the plasma current
@@
(b) $\kappa_{2,\rho}$ umbilic coil
@@
The effect of the umbilic coil can be decomposed as a rigid shift of the whole equilibrium and a small deformation of the boundary. Since the rigid shift (Sengupta et al., 2024) and the size of the ridge is small compared to the minor radius, it does not compromise omnigenity (as seen later in figure 11(a)). For a perfectly sharp umbilic edge, the magnetic field line must coincide with the edge, but even when the ridge lacks perfect sharpness, the magnetic field line stays close to it due to its high curvature.
[…]
@@
We observe the presence of a stochastic layer outside the plasma boundary, similar to heliotron/torsatron configurations. Since the plasma and the umbilic coil are carrying currents in opposite directions, the poloidal field cannot vanish between them, which is why we do not observe an X-point. Therefore, the umbilic coil can create and vary the curvature of the ridge without creating an varying the current in the umbilic coil without creating an X point. This also affects the fraction of the rotational transform generated by the umbilic coils. These details are provided in figure 6.
@@
We can see that the omnigenity degrades as we increase $I_{\mathrm{umbilic}}$ but the neoclassical losses are still acceptable $\epsilon_{\mathrm{eff}}^{3/2}<0.002$. To control instabilities and transport due to MHD modes, we may need to generate a higher $\iota_{\mathrm{shaping}}$ while ensuring omnigenity. One way to explore multiple such coil and equilibrium configurations would be to perform a single-stage optimization, which simultaneously changes the plasma boundary and coil shapes. This is beyond the scope of the current study.
5.2 Umbilic coil current in the same direction as the plasma current
In a manner similar to the previous section, we create a setup where we modify the HBT-EP experiment and add a $m=n=1$ ridge to the boundary. However, we do this while ensuring that the toroidal component of the current in the umbilic coil $I_{\mathrm{umbilic}}=-6.1kA$ is in the same direction as the plasma current $I=-35.2kA$. The coilset and free-boundary solution are presented in figure 12.
@@ 5.2 Umbilic coil current in the same direction as the plasma current
(b) $\kappa_{2,\rho}$ umbilic coil
@@
Figure 12: Coilset and free-boundary solution for the case with the coils are closer to the plasma and the current in the umbilic coil points in the same direction as the plasma current and since no stage-two optimization was performed the second-stage objective $\mathcal{F}_{\mathrm{stage-two}}$ has a large value.
Due to the proximity of the coil to the plasma, the maximum $\lvert\boldsymbol{B}\cdot\hat{\boldsymbol{n}}\rvert$ and pressure balance errors were larger than the acceptable normalized value (of $1\%$), implying the boundary shape will differ from the one shown in figure 12.
[…]
Since the umbilic coil can reduce the poloidal field significantly, we hypothesize that it should be possible to optimize for the coil position and plasma boundary shape so that there is a region between the plasma and the coil that behaves like an X point. This could potentially convert the initially limited HBT-EP into a diverted configuration.
@@ 6 Conclusions and future work
This paper explored the physical properties of umbilic stellarators. In section 3, we defined the umbilic parametrization, approximated it using DESC and solved the ideal MHD equation inside the enclosed volume. In section 4, we developed vacuum and finite-$\beta$ stellarators by simultaneously optimizing the umbilic edge and the boundary shape using the DESC optimization suite. We designed coils for the vacuum case, and by tracing the field lines in the edge, showed how the high-curvature edges could be a favorable location to place divertors. We found that the optimized omnigenous umbilic designs do not possess a specific helicity regardless of the low-ripple transport, which could be a way to generate piecewise omnigenous equilibria. As a final demonstration, in section 5, we propose an experiment to modify the Columbia HBT-EP experiment to an umbilic design by adding a $m=1,n=1$ helical coil. Using the DESC free boundary solver, we demonstrate that the umbilic coils are able to cause the originally circular plasma to bulge and form a high-curvature ridge while maintaining omnigenity. By changing the direction of the current in the umbilic coil, we may be able to get an X-point, thereby converting the limited HBT-EP tokamak into a diverted stellarator.
@@
The technique of simultaneously optimizing a curve and a surface to control the magnitude and position of a sharp edge along a surface extends beyond umbilic stellarators. One can use this method to place and design ridges or regions of high curvature which may be useful for non-resonant divertors (Bader et al., 2018b), as well as these ridges creating preferential pathways for the field lines to exit the plasma (Bader et al., 2017). For an island divertor concept, the rotational transform on the boundary can be fine-tuned using umbilic coils, which could be used to reduce the sensitivity of X-points in the island divertor structure due to changes in the plasma boundary rotational transform. This idea is similar to that of divertor coils in tokamaks and would lead to a predictable strike point pattern.
[…]
@@
Acknowledgements One of the authors, R. G. enjoyed discussions with M. Zarnstroff, J. Levesque, J. L. Velasco, B. Jang, and gratefully acknowledges the help from S. Lazerson installing the FIELDLINES code. This work is funded through the SciDAC program by the US Department of Energy, Office of Fusion Energy Science, and Office of Advanced Scientific Computing Research under contract No. DE-AC02-09CH11466, the Hidden Symmetries grant from the Simons Foundation/SFARI (560651), and Microsoft Azure Cloud Computing mini-grant awarded by the Center for Statistics and Machine Learning (CSML), Princeton University. This research also used the Della cluster at Princeton University.
@@
Data availability statement The data and analysis files are freely available in Gaur (2025).
@@ Appendix C Piecewise omnigenity in umbilic configurations
[…]
@@
For a perfect, non-piecewise omnigenous stellarator $\partial_{\alpha}(\mathcal{J}_{\parallel})=0$, which means that $\mathcal{J}_{\parallel}$ is continuous and constant in $\alpha$ for almost every $1/\lambda$. Moreover, the distinct $\mathcal{J}_{\parallel}$ can be written as a linear sum such that the total $\mathcal{J}_{\parallel}$ for a value of the inverse pitch vanishes. However, for the umbilic configurations in this paper, especially the US131 vacuum stellarator, despite a low effective ripple (good omnigenity), we clearly see that $\mathcal{J}_{\parallel}$ is discontinuous in $\alpha$ for most of the inverse pitch values and $\tilde{\mathcal{J}}_{\parallel,\mathrm{I}}\approx\tilde{\mathcal{J}}_{\parallel,\mathrm{II}}\approx 0.5\tilde{\mathcal{J}}_{\parallel,\mathrm{III}}$. This supports our claim that the umbilic stellarator US131 is piecewise omnigenous, especially at large values of the inverse pitch. US252 seems to be piecewise omnigenous but the structure of $\tilde{\mathcal{J}_{\parallel}}$ is more complicated than the one discussed by Velasco et al. (2024).
@@ Appendix D Details of the HBT-EP equilibrium
The original HBT-EP equilibrium is a shifted-circle equilibrium with a major radius $R_{0}=0.94m$, minor radius $a_{0}=0.125m$ (aspect ratio $A=7.52$) with a $\beta=0.5\%$ and a plasma current $I=-35.2kA$. The pressure and rotational transform profiles are given in figure 14. The initial equilibrium was tested to be stable against $n=1$ kink modes using the DCON-2D (Glasser, 2016) stability code.

### v2

@@ 5 Umbilic coil design for a finite-$\beta$ equilibrium
As we have seen in section 4.1, creating modular coils for vacuum umbilic equilibria is a difficult task due to the sharp curvature regions. Therefore, in this section, we present an alternate strategy that uses umbilic coils to create a US111 finite-$\beta$ equilibrium. Instead of modifying modular coils, we shall use an umbilic coil, a low-current, helically wound coil, similar to a divertor coil in tokamaks, which can be used to create an umbilic edge. Unlike tokamak divertor coils, an umbilic coil is non-axisymmetric and can have a fractional helicity. In the following sections, we will discuss two cases, categorized by the direction of the current in the umbilic coils with respect to the toroidal plasma current.
5.1 US111 with an umbilic coil current opposite to the plasma current (counter-$I$ case)
@@
(b) $\kappa_{2,\rho}$ & umbilic coil
@@
The effect of the umbilic coil can be decomposed as a rigid shift of the whole equilibrium and a small deformation of the boundary. Since the rigid shift (Sengupta et al., 2024) and the size of the ridge is small compared to the minor radius, it does not compromise omnigenity (as seen later in figure 12(a)). For a perfectly sharp umbilic edge, the magnetic field line must coincide with the edge, but even when the ridge lacks perfect sharpness, the magnetic field line stays close to it due to its high curvature. The curvature along the umbilic ridge $\kappa_{1,\rho}\in[-15,-32]m^{-1}$ whereas the average curvature $\mathrm{avg}(\kappa_{2,\rho})=-7m^{-1}$.
The TF coils each carry a current of $105\,\text{kA}$, the PF coils carry $29.8\,\text{kA}$, and the umbilic coil has a current $I_{\mathrm{umbilic}}=2.1\,\text{kA}$. But due to the proximity of the umbilic coil compared to TF and PF coils, it can still affect the plasma near the boundary significantly. As we can see, the umbilic coil is able to create a $m=1,n=1$ high curvature ridge. The ridge formation is entirely attributed to the umbilic coil, as all other coils are axisymmetric. The current flowing through the coil is in the opposite direction to the plasma current. However, due to the 3D nature of this setup, the plasma still bulges toward the coil.
@@
We observe the presence of a stochastic layer outside the plasma boundary, similar to heliotron/torsatron configurations. Since the plasma and the umbilic coil are carrying currents in opposite directions, the poloidal field cannot vanish between them, which is why we do not observe an X-point. Therefore, the umbilic coil can create and vary the curvature of the ridge without creating and varying the current in the umbilic coil without creating an X point. This also affects the fraction of the rotational transform generated by the umbilic coils. These details are provided in figure 6.
@@
We can see that the omnigenity degrades as we increase $I_{\mathrm{umbilic}}$ but the neoclassical losses are still acceptable $\epsilon_{\mathrm{eff}}^{3/2}<0.001$. To control instabilities and transport due to MHD modes, we may need to generate a higher $\iota_{\mathrm{shaping}}$ while ensuring omnigenity. One way to explore multiple such coil and equilibrium configurations would be to perform a single-stage optimization, which simultaneously changes the plasma boundary and coil shapes. This is beyond the scope of the current study.
5.2 US111 with an umbilic coil current in the same direction as the plasma current (co-$I$ case)
In a manner similar to the previous section, we create a setup where we modify the HBT-EP experiment and add a $m=n=1$ ridge to the boundary. However, we do this while ensuring that the toroidal component of the current in the umbilic coil $I_{\mathrm{umbilic}}$ is in the same direction as the plasma current I=35.2kAI=-35.2kA. To find the umbilic coil shape, coil current, and the corresponding plasma shape that minimizes the normal field error $\boldsymbol{B}\cdot\hat{\boldsymbol{n}}$ and the total pressure across the plasma boundary, we combine the objectives defined in (14) and (16) in a process known as single-stage optimization. We solve
$\min(\mathcal{F}_{\mathrm{stage-one}}+\mathcal{F}_{\mathrm{stage-two}}),\quad\textrm{s.t.}\,\boldsymbol{\nabla}\left(\mu_{0}p+\frac{B^{2}}{2}\right)-\boldsymbol{B}\cdot\boldsymbol{\nabla}\boldsymbol{B}=0,\,\psi_{\mathrm{b}}=\psi_{\mathrm{b}0},p=p_{0}(\psi),I=I_{0}(\psi)$ (17)
where $\psi_{\mathrm{b}0}$ is the user-provided enclosed toroidal flux by the boundary, $p_{0}(\psi)$ and $I_{0}(\psi)$ are the pressure and toroidal current profiles. The optimizable set of parameters
[…]
@@ 5.2 US111 with an umbilic coil current in the same direction as the plasma current (co-$I$ case)
(b) $\kappa_{2,\rho}$ & umbilic coil
@@
Figure 13: Coilset and free-boundary solution for the case where the current in the umbilic coil is in the same direction as the plasma current. The current carried by the umbilic coil $I_{\mathrm{umbilic}}=-3.1kA$. All coils but the umbilic coil are axisymmetric.
After optimization, the normalized $\mathrm{max}(\boldsymbol{B}\cdot\hat{\boldsymbol{n}})<0.5\%$ and the total pressure difference is $0.1\%$. Because we allow the boundary shape to change, we have less control over the curvature along the umbilic edge, so there is a lot of variation. The curvature along the edge lies in the range $\kappa_{2,\rho}\in[-25,-13]\,m^{-1}$ with the average curvature $\mathrm{avg}(\kappa_{2,\rho})=-8m^{-1}$. The neoclassical transport coefficient ${\epsilon_{\mathrm{eff}}}^{3/2}<0.001$ implies that the magnetic field is still omnigenous. Finally, we calculate the rotational transform profile of the optimized equilibrium to quantify the effect of the umbilic coil in figure 14.
Figure 14: The total rotational transform profiles and its contribution from shaping and plasma current. The non-axisymmetric shaping is purely a consequence of the umbilic coil which contributes to approximately $5\%$ of the total rotational transform in the edge.
The effect on the rotational transform profile due to the umbilic coil rises sharply as we move towards the edge. Therefore it is possible that outside the plasma boundary the umbilic coil is able to affect the total magnetic field in such a way that it forms an X-point.
@@ 6 Conclusions and future work
This paper explored the physical properties of umbilic stellarators. In section 3, we defined the umbilic parametrization, approximated it using DESC and solved the ideal MHD equation inside the enclosed volume. In section 4, we developed vacuum and finite-$\beta$ stellarators by simultaneously optimizing the umbilic edge and the boundary shape using the DESC optimization suite. We designed coils for the vacuum case, and traced fieldlines in the edge of the US131 stellarator before and after adding a current source at the magnetic axis, to simulate a plasma fluctuation. Even though we do not see a perfectly sharp X-point (X-line) and well-defined fieldline structure like a tokamak, the high-curvature edges could be a favorable location to place divertors and are robust to internal plasma fluctuations. We found that the optimized omnigenous umbilic designs do not possess a specific helicity regardless of the low-ripple transport, which could be a way to generate piecewise omnigenous equilibria. As a final demonstration, in section 5, we propose an experiment to modify the Columbia HBT-EP experiment to an umbilic design by adding a $m=1,n=1$ helical coil. Using the DESC free boundary solver, we demonstrate that the umbilic coils are able to cause the originally circular plasma to bulge and form a high-curvature ridge while maintaining omnigenity. By changing the direction of the current in the umbilic coil, we may be able to get an X-point, thereby converting the limited HBT-EP tokamak into a diverted stellarator.
@@
The technique of simultaneously optimizing a curve and a surface to control the magnitude and position of a sharp edge along a surface extends beyond umbilic stellarators. One can use this method to place and design ridges or regions of high curvature which may be useful for non-resonant divertors (Bader et al., 2018b), as well as these ridges creating preferential pathways for the field lines to exit the plasma (Bader et al., 2017). This can be achieved by multiplying a custom function $g(\theta)$ to the objective $f_{\mathrm{umbilic}}$ such that $g(\theta)=1\quad\forall\quad\theta\in\bigcup_{i}[\theta_{i0},\theta_{i1}]$ and $g(\theta)=0$ elsewhere along the umbilic curve, with $\theta_{i0},\theta_{i1}$ being optimizable parameters describing the lower and upper limits of regions where low curvature is penalized. This will create a set of continuous but disconnected ridges, just as we observe in non-resonant divertor concepts. This functionality is available in DESC. For an island divertor concept, the rotational transform on the boundary can be fine-tuned using umbilic coils, which could be used to reduce the sensitivity of X-points in the island divertor structure due to changes in the plasma boundary rotational transform. This idea is similar to that of divertor coils in tokamaks and would lead to a predictable strike point pattern.
Future work would involve using multiple umbilic coils to ensure plasma stability against low-$n$ modes ($n=0$ family) and increase the curvature of the ridge, similar to the work on tokamak separatrix by Webster (2009), understanding the strike point pattern and field line structure outside the plasma boundary or designing umbilic coils for a case similar to the finite-$\beta$ $\rm{US}252$ stellarator in section 4.4. It is also important to include MHD and kinetic stability metrics in the optimization process, especially near the boundary as has been done for tokamaks by Zheng et al. (2025); Bishop (1986) as stability properties can be sensitive close to the high-curvature plasma boundary.
Acknowledgements One of the authors, R. G. enjoyed discussions with M. Zarnstroff, J. Levesque, R. Davies, J. L. Velasco, B. Jang, and gratefully acknowledges the help from S. Lazerson installing the FIELDLINES code. We also express our gratitude to the referees for their constructive feedback that has improved the overall quality of this paper.
@@
Funding This work is funded through the SciDAC program by the US Department of Energy, Office of Fusion Energy Science, and Office of Advanced Scientific Computing Research under contract No. DE-AC02-09CH11466, the Hidden Symmetries grant from the Simons Foundation/SFARI (560651), and Microsoft Azure Cloud Computing mini-grant awarded by the Center for Statistics and Machine Learning (CSML), Princeton University. This research also used the Della cluster at Princeton University.
@@
Data availability statement All the data, analysis, and post-processing files are freely available in Gaur (2025).
Declaration of interests Matt Landreman is a consultant for Type One Energy
@@ Appendix C Piecewise omnigenity in umbilic configurations
In this appendix, we will plot the normalized second adiabatic invariant $\mathcal{J}_{\parallel}=\int_{l_{1}}^{l_{2}}dl\,v_{\parallel}/(2\pi R_{0})$ as a function of the field line label $\alpha$ and the inverse pitch angle $1/\lambda\equiv E/\mu$. We calculate $\mathcal{J}_{\parallel}(\alpha,1/\lambda)$ on a single flux surface on which the neoclassical transport coefficient $\epsilon_{\mathrm{eff}}^{3/2}\leq 0.001$. This is done for both the US131 vacuum and US252 finite-beta equilibria presented in sections 4.1 and 4.4, respectively. To highlight their difference, we also plot $\mathcal{J}_{\parallel}$ for a poloidally omnigenous (or QI) $n_{\rm{FP}}=1$ vacuum equilibrium taken from the DESC omnigenity database (Gaur et al., 2023). Comparison is shown in figure 15.
@@
For a perfect, non-piecewise omnigenous stellarator $\partial_{\alpha}(\mathcal{J}_{\parallel})=0$, implying that $\mathcal{J}_{\parallel}$ is continuous and constant in $\alpha$ for almost every $1/\lambda$. Moreover, the distinct $\mathcal{J}_{\parallel}$ can be written as a linear sum such that the total $\mathcal{J}_{\parallel}$ for a value of the inverse pitch vanishes. However, for the umbilic configurations in this paper, especially the US131 vacuum stellarator, despite a low effective ripple (good omnigenity), we clearly see that $\mathcal{J}_{\parallel}$ is discontinuous in $\alpha$ for most of the inverse pitch values and $\tilde{\mathcal{J}}_{\parallel,\mathrm{I}}\approx\tilde{\mathcal{J}}_{\parallel,\mathrm{II}}\approx 0.5\tilde{\mathcal{J}}_{\parallel,\mathrm{III}}$. This supports our claim that the umbilic stellarator US131 is piecewise omnigenous, especially at large values of the inverse pitch. US252 seems to be piecewise omnigenous but the structure of $\tilde{\mathcal{J}_{\parallel}}$ is more complicated than the one discussed by Velasco et al. (2024).
@@ Appendix D Details of the HBT-EP equilibrium
The original HBT-EP equilibrium is a shifted-circle equilibrium with a major radius $R_{0}=0.94m$, minor radius $a_{0}=0.125m$ (aspect ratio $A=7.52$) with a $\beta=0.5\%$ and a plasma current I=35.2kAI=-35.2kA. The pressure and rotational transform profiles are given in figure 16. The initial equilibrium was tested to be stable against $n=1$ kink modes using the DCON-2D (Glasser, 2016) stability code.